OpenStudy (anonymous):

Question 3, @satellite73 this one has 4 parts so it might be a little longer, if you need to go or want to finish tomorrow just let me know i wont mind (: thanks for all your help today btw

OpenStudy (anonymous):

lets take a look

OpenStudy (anonymous):

first part is real easy replace \(s\) by \(96\) and \(v\) by \(80\) and get \[H(t)=-16t^2+80t+96\]

OpenStudy (anonymous):

second part is also easy the vertex is where the first coordinate is \(-\frac{b}{2a}\) which in our case is \(-\frac{80}{2\times (-16)}=\frac{5}{2}\)

OpenStudy (anonymous):

or if you prefer decimals \(2.5\)

OpenStudy (anonymous):

then the maximum height will be \[H(2.5)=-16\times (2.5)^2+80\times (2.5)+96\] which you compute using a calculator

OpenStudy (anonymous):

i get this http://www.wolframalpha.com/input/?i=-16%282.5%29^2%2B80*2.5%2B96

OpenStudy (anonymous):

I am sorry had to go to the bathroom give me a second to catch up (:

OpenStudy (anonymous):

take your time

OpenStudy (anonymous):

Okay, read it all and wow yeah part A was easy lol how did i not see that

OpenStudy (anonymous):

part B i hope you get, if didn't leave out any steps

OpenStudy (anonymous):

yes i understood that too dont worry (:

OpenStudy (anonymous):

plus i used a calculator, so i know the answer is right

OpenStudy (anonymous):

C want you to do \[-16t^2+80t+96=31+32.2t\]

OpenStudy (anonymous):

it says use a table, i would use this http://www.wolframalpha.com/input/?i=-16t^2%2B80t%2B96%3D31%2B32.2t

OpenStudy (anonymous):

ignore the negative solution because time does not go backwards round and say it is at about \(t=4\)

OpenStudy (anonymous):

or you could just plug in 4 and say it is close

OpenStudy (anonymous):

okay yeah thanks, anyway if you were to round the answer it would be a 4 right? (:

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and finally D is also easy it goes up until time is \(2.5\) because that was the first coordinate of the vertex since \(4>2.5\) it is on the way DOWN

OpenStudy (anonymous):

Okay, thank you soooo much (: <3

OpenStudy (anonymous):

yw you owe me a beer good luck!

OpenStudy (anonymous):

when im old enough to get one then you will get your beer :P

OpenStudy (anonymous):

oops i will settle for a root beer

OpenStudy (anonymous):

lol xD well anyway thanks a bunch this really means a lot considering i have eoc exam on monday and now i can take it since you helped me with this, all i need to do now is go back and write down the answers, so yeah thanks xoxo

OpenStudy (anonymous):

yw good luck don't study too late

OpenStudy (anonymous):

thank you and haha will do (: